A horizontal pipe contains water at a pressure of 150 kPa flowing with a speed of 1.8 m/s . The pipe narrows to one-half its original diameter.What is the speed of the water?What is the pressure of the water?
Answer #1

Here,

v1 = 1.8 m/s

P1 = 150 kPa

when the diameter becomes one half , the area of cross section will become 1/4 times

and the speed become 4 times

speed of water , v2 = 4 * v1

speed of water , v2 = 4 * 1.8

speed of water , v2 = 7.2 m/s

the speed of water is 7.2 m/s

Now, for the pressure P2

P1 + 0.50 p * v1^2 = P2 + 0.50 * p * v2^2

150 * 10^3 + 0.50 * 1000 * 1.8^2 = P2 + 0.50 * 1000 * 7.2^2

P2 = 125700 Pa = 125.7 kPa

the pressure P2 is 125.7 kPa